A spherical shell contains three charged objects T
Answer #1

According to Gauss's law net flux through a surface is \oint{\vec{E}\cdot\,d\vec{S}}=\frac{Q_{enc}}{\epsilon_0}

Here net flux is \phi=-953\,N.m^2/C

enclosed charge is Q_{enc}=-18.0\times10^{-9}+37.0\times10^{-9}+q

Q_{enc}=\phi\epsilon_0=-18.0\times10^{-9}+37.0\times10^{-9}+q=(-953)\times(8.854\times10^{-12})

-18.0\times10^{-9}+37.0\times10^{-9}+q=-8.44\times10^{-9}\,C

charge on the third object is q=-27.44\times10^{-9}\,C